OpenStudy (anonymous):

Help! Can anyone walk me through solving this for both variables? 8x+12y=24

OpenStudy (dtan5457):

Isolate each variable at a time 8x+12y=24 8x=24-12y 12y=24-8x divide x and y accordingly